OpenStudy (anonymous):

What are the eigenvectors of \[\left[\begin{matrix}7 & -\sqrt{3} \\ -\sqrt{3} & 5\end{matrix}\right] \]

OpenStudy (anonymous):

\[\left[\begin{matrix}7- \lambda & - \sqrt{3} \\ - \sqrt{3} & 5- \lambda \end{matrix}\right]\] Solve for Eigenvalues \( \lambda \) first, then you can discuss the eigenvectors.

OpenStudy (anonymous):

The eigenvalues are 8 and 4

OpenStudy (anonymous):

What would I do now?

OpenStudy (anonymous):

Exactly, OpenStudy is kicking me pretty often today so bear with me. You can substitute back the Eigenvalues. \( \lambda_1=4 \) You get the following system of equations: \[3x - \sqrt{3}y=0 \\- \sqrt{3}x + y=0 \]

OpenStudy (anonymous):

If you multiply the lower equation by \( - \sqrt{3} \) you will see that they are identical. So you can only solve one of these systems.

OpenStudy (anonymous):

Yeah, having the same problem with os. Alright, so I found the eigenvectors as: \[\left(\begin{matrix}-\sqrt{3} \\1\end{matrix}\right) \ and \ \left(\begin{matrix}1/\sqrt{3} \\ 1\end{matrix}\right)\] are these correct?

OpenStudy (anonymous):

If you look at the first equation you have \[ 3x-\sqrt{3}y=0 \] This means that: \[\vec{n}=\left(\begin{matrix}3 \\ -\sqrt{3}\end{matrix}\right)\] is a normal vector of this equation, so if you want the eigenvector you just switch the columns and change one sign, so you should get: \[\left(\begin{matrix}\sqrt{3} \\ 3\end{matrix}\right)\] as an Eigenvector.

OpenStudy (anonymous):

Wouldn't the equations be: -x-sqrt(3)=0 3x-sqrt(3)y=0

OpenStudy (anonymous):

Wait, seems you have that. Both equations work.

OpenStudy (anonymous):

Well if it works the way you did it stick to it, the general outcome is that one equation becomes redundant, because it's just a multiple of the other equation, so you can forget about it and solve for Eigenvectors using one equation.

OpenStudy (anonymous):

Alright, thanks

OpenStudy (anonymous):

welcome